OpenStudy (anonymous):

Write the expression as a single natural logarithm. 2 ln 8 + 2 ln y

OpenStudy (callisto):

Use the following properties:\[b \ln a = \ln a^b\]\[\ln m + \ln n = \ln (mn)\]